# Tinniam V Ganesh

### 15 Apr 2026 v1

@@ Abstract
This paper develops a unified Markov Decision Process (MDP) framework for optimising two recurring in-match decisions in T20 cricket viz. batting order selection and bowling plan assignment, directly in terms of win and defend probability rather than expected runs. A three-phase player profile engine (Powerplay, Middle, Death) with James–Stein shrinkage is estimated from 1,161 IPL ball-by-ball records (2008–2025). Win/defend probabilities are evaluated by vectorised Monte Carlo simulation over N=50, [^0] N=50{,}000 innings trajectories. Batting orders are searched by exhaustive enumeration. Bowling plans are computed by simulated annealing over the remaining quota with the constraint that the same bowler cannot bowl consecutive overs. Applied to two 2026 IPL matches, the optimal batting order improves Mumbai Indians’ win probability by 4.1 pp (52.4% → 56.5%), and the optimal Gujarat Titans bowling plan improves defend probability by 5.2 pp (39.1% → 44.3%). In both cases the source of suboptimality is phase-agnostic deployment of players.
@@ 1. Introduction
    Simulation-based search, namely exhaustive enumeration for batting order and simulated annealing for bowling plans, that directly optimises win/defend probability over the stochastic innings model, validated at N=50, [^0] N=50{,}000 simulations per configuration (Monte Carlo SE $\approx 0.22\%$).
@@ 4.5 James–Stein Shrinkage for Phase-Sparse Players
Many IPL players have limited exposure in specific phase-role combinations: a seam bowler who has rarely bowled Middle overs, a spinner who has never bowled Death overs, or a new franchise addition with fewer than two full IPL seasons. For such players, the Laplace-smoothed individual estimate $\hat{\mathbf{p}}_{i}^{\phi}$ is a high-variance estimate from few observations. We apply a James–Stein shrinkage estimator (James and Stein, 1961), which blends each player’s individual estimate toward the population-average profile for that phase.

### 16 Apr 2026 v2

@@ Abstract
This paper develops a unified Markov Decision Process (MDP) framework for optimising two recurring in-match decisions in T20 cricket, namely batting order selection and bowling plan assignment, directly in terms of win and defend probability rather than expected runs. A three-phase player profile engine (Powerplay, Middle, Death) with James–Stein shrinkage (a technique that blends a player’s individual statistics toward the league average when their phase-specific data is sparse) is estimated from 1,161 IPL ball-by-ball records (2008–2025). Win/defend probabilities are evaluated using vectorised Monte Carlo simulation over N=50, [^0] N=50{,}000 innings trajectories.
Batting orders are evaluated by comparing all feasible arrangements of the remaining players and selecting the one that maximises win probability. Bowling plans are optimised through a guided search over possible over assignments, progressively improving the allocation while respecting constraints such as the prohibition on consecutive overs by the same bowler.
Applied to two 2026 IPL matches, the optimal batting order improves Mumbai Indians’ win probability by 4.1 percentage points (52.4% to 56.5%), and the optimal Gujarat Titans bowling plan improves defend probability by 5.2 percentage points (39.1% to 44.3%). In both cases, the observed sub-optimality is consistent with phase-agnostic deployment: decisions that appear reasonable under aggregate metrics are shown to be costly when phase-specific profiles are applied.
@@ 1. Introduction
    Simulation-based search: for batting, all feasible orderings of the remaining players are evaluated and the one maximising win probability is selected; for bowling, a guided search progressively improves the over assignment while respecting constraints such as the prohibition on consecutive overs by the same bowler. Both directly optimise win/defend probability over the stochastic innings model, validated at N=50, [^0] N=50{,}000 simulations per configuration (Monte Carlo SE $\approx 0.22\%$).
@@ 4.5 James–Stein Shrinkage for Phase-Sparse Players
Many IPL players have limited exposure in specific phase-role combinations: a seam bowler who has rarely bowled Middle overs, a spinner who has never bowled Death overs, or a new franchise addition with fewer than two full IPL seasons. For such players, the Laplace-smoothed individual estimate $\hat{\mathbf{p}}_{i}^{\phi}$ is a high-variance estimate from few observations. James–Stein shrinkage addresses this by blending each player’s own phase-specific estimate toward the population average for that phase, so that the less data a player has, the more their profile is pulled toward the league mean. We apply a James–Stein shrinkage estimator (James and Stein, 1961) to implement this.